Show that if u1 , u2, u3, u4 are linearly
independent vectors in a vector space V over
a field K, then u 1+ u2, u3— u4, u4+ u1 are
also linearly independent.
1
Expert's answer
2020-11-19T17:10:51-0500

Given "V" is a vector space over the field "K" .

"u_1,u_2,u_3,u_4" are linearly independent"(L.I)" vectors in "V" .

Now, let for all "a,b,c\in K" , consider the linear combination


"a(u_1+u_2)+b(u_3-u_4)+c(u_4+u_1)=0\\
\implies (a+c)u_1+au_2+bu_3+(c-b)u_4=0\\
\implies (a+c)=a=b=(c-b)=0\hspace{1cm}(\because u_1,u_2,u_3,u_4 \,L.I)\\
\implies a=b=c=0"

Thus,"u_1+u_2,u_3-u_4,u_4+u_1" are "L.I" .